OpenStudy (anonymous):

Which ratio is equivalent to 4 : 28? medals!!!!!

OpenStudy (anonymous):

1 : 7 28/4 = 7

OpenStudy (anonymous):

what raito is proportional to 10/24

OpenStudy (anonymous):

1 : 2.4

OpenStudy (anonymous):

35/84 5/6 5/8 20/36

OpenStudy (anonymous):

1 : 2.4 1: 1.2 1 : 1.6 1 : 1.8

OpenStudy (anonymous):

hun thoes are the answers to the question

OpenStudy (anonymous):

what raito is proportional to 10/24 35/84 5/6 5/8 20/36

OpenStudy (anonymous):

look closer and I already answered it lol

OpenStudy (anonymous):

10/24 = 35/84

OpenStudy (anonymous):

i am so confused..!! lol